OpenStudy (anonymous):

does integral ( integral ( (x-y)/(x^2+y*2) ) ) converges or not? (it is double integral) I have this exercise I don't know how to do it I even don't know what does that means , so please help

OpenStudy (anonymous):

The integral is over what set of R^2 ?

OpenStudy (anonymous):

oh sorry forgot so x+y>1 , x>0 , y>0

OpenStudy (anonymous):

the denominator ..x^2+y^2 is it power or multiplication

OpenStudy (anonymous):

@onikadze

OpenStudy (anonymous):

it is power

OpenStudy (anonymous):

hmm... kk

OpenStudy (anonymous):

@onikadze whts the ans given i am doing same sums soo.... can give the ans

OpenStudy (anonymous):

shall I tell you answer?

OpenStudy (anonymous):

okay the answer is that it is not converged

OpenStudy (anonymous):

here is the way how it was solved 35 th exersise

OpenStudy (anonymous):

@shkrina

OpenStudy (anonymous):

@onikadze i am sorry i have just stared that it too deep ... they can help u @amistre64 @mathstudent55 @Luigi0210

OpenStudy (anonymous):

okay thanks anyway

OpenStudy (anonymous):

@cwrw238

OpenStudy (luigi0210):

@Jhannybean

OpenStudy (anonymous):

\[\int\int_{\Omega}\frac{x-y}{x^2+y^2}~dx~dy\] with \(\Omega:=\left\{(x,y)~|~x+y>1,~x>0,~y>0\right\}\). Here's a sketch of the region \(\Omega\):  Converting to polar coordinates, your integral becomes \[\int\int_{\Omega}\frac{r\cos\phi-r\sin\phi}{r^2}~dr~d\phi\]